### v1

# Improved Supervised Fine-Tuning for Large Language Models to Mitigate Catastrophic Forgetting

Abstract
Supervised Fine-Tuning (SFT), while enhancing large language models(LLMs)’ instruction-following capabilities and domain-specific task adaptability, often diminishes their general capabilities. Moreover, due to the inaccessibility of original pre-training data, catastrophic forgetting tends to be exacerbated when third-party practitioners implement SFT on open-sourced models. To address this challenge, we propose a novel, more cost-effective SFT method which could effectively reduce the risk of catastrophic forgetting without access to original SFT data. Our approach begins by reconstructing the likely SFT instruction distribution of the base model, followed by a multi-model screening process to select optimal data, which is then mixed with new data for SFT. Experimental results demonstrate that our method preserves generalization capabilities in general domains while improving task-specific performance.
Improved Supervised Fine-Tuning for Large Language Models to Mitigate Catastrophic Forgetting
Fei Ding, Baiqiao Wang Correspondence: dingfei@email.ncu.edu.cn
1 Introduction
Large Language Models (LLMs) have demonstrated remarkable proficiency in understanding and generating human language, leading to their widespread adoption across diverse applications. Meanwhile, recent advancements have produced powerful, publicly accessible base LLMs, including LLaMA  Touvron et al. (2023a, b), Falcon Penedo et al. (2023), and Mistral  Jiang et al. (2023). To optimize performance in specialized domains such as medicine, law, and science, fine-tuning these general-purpose LLMs on domain-specific datasets is often necessary. However, this process introduces a critical challenge: catastrophic forgetting(McCloskey and Cohen, 1989), where models lose previously acquired general knowledge, thereby limiting their broader applicability and reusability across domains.
To mitigate catastrophic forgetting, a series of rehearsal-based methods have been proposed. However, traditional rehearsal-based methods mainly focus on the continual learning process and relies heavily on training data.To mitigate the degradation of general capabilities, the specific way is to blend general SFT data with domain-specific data, while mainstream open-source models like LlamA and Qwen do not disclose their original SFT datasets. Current practices often involve mixing public SFT datasets with custom data, but these public datasets frequently diverge significantly from the base model’s original SFT data distribution, exacerbating SFT performance degradation. Therefore, better SFT requires an alternative approach to deal with models that do not disclose their original SFT data. To address this challenge, we propose a novel, cost-effective rehearsal-based method that reconstructs the hidden instruction data distribution of base models and then generate synthetic SFT data through multi-response generation and filter for SFT. Furthermore, our method is compatible with existing continual learning approaches, enabling more effective sequential fine-tuning.
Our contributions are as follows:
  • We propose a method that could effectively reconstruct the likely hidden instruction distribution of open-source models by analyzing their generative behavior, guiding the creation of general-purpose SFT datasets to mitigate the degradation of general capabilities.
  • We introduce a Multi-Response Generation and Filter mechanism, which optimizes instruction-response pairs via mutual scoring across three models, significantly improving response quality and SFT efficacy.
  • Experiment results demonstrate that our SFT method can significantly reduce the decline in the LLM’s general capabilities after fine-tuning compared to directly mixing third-party datasets.
2 Related Work
2.1 Catastrophic Forgetting
Continual learning is essential for models that must adapt to dynamic environments, assimilating information from a continuous data stream while retaining previously learned knowledge. A critical obstacle in this domain is catastrophic forgetting, which refers to the pronounced degradation in performance on earlier tasks when the model is exposed to novel data. As the model adjusts its parameters to incorporate new information, it inadvertently overwrites previously acquired knowledge, thereby diminishing its performance on prior tasks. To address catastrophic forgetting, existing continual learning approaches can be broadly categorized into four classes: Replay-Based Shin et al. (2017); Ren et al. (2024), Regularization-Based Mi et al. (2020), Gradient-Based Lee et al. (2021), and Architecture-Based Geng et al. (2021) approaches.
2.2 Data Rehearsal
Robins ROBINS (1995) introduced the concept of rehearsal shortly after the identification of catastrophic forgetting. Rehearsal aims to mitigate forgetting by re-exposing the model to data or representations from previously learned tasks. Common rehearsal methods include Experience Replay Rolnick et al. (2019), storing subsets of past task examples and periodically retraining the model on a mixture of these historical examples and new data, and Generative Replay Shin et al. (2017), using generative models to create synthetic examples of prior tasks for integration into new task training. A recent innovation is Self-Synthesized Rehearsal (SSR) Huang et al. (2024), which continuously leverages LLM checkpoints from prior stages to synthesize and refine input instances, blending them with current-stage data.
3 Methods
Our framework comprises two stages: (1) reconstructing the original SFT data distribution from the base model, and (2) blending the reconstructed data with new domain-specific data for training.
3.1 Reconstructing General-Purpose SFT Data from LLMs
The process of recovering the instruction distribution involves three automated steps: instruction generation, response generation, and high-quality response filtering.
Step 1: Instruction Generation. The objective of this step is to approximate the distribution of instructions used in the model’s original SFT phase. Given an open-weight aligned large language model (e.g., Llama-3-70B-Instruct), we design a pre-query template in the format of the predefined instruction template. We input the prompt "<|start_header_id|>user<|end_header_id|>" into the large model (Llama-3-70B-Instruct), which generates a single instruction in response. By repeating this process 100,000 times, we obtain a total of 100,000 instructions, which collectively represent the current instruction distribution of the large model.
Step 2: Multi-Response Generation. The objective of this step is to generate responses to the instructions obtained in Step 1.
We send these instructions to Llama-3-70B-Instruct and two additional powerful large language models (such as gpt4 and Qwen2.5-72B-Instruct). For each instruction, each model generates three responses, resulting in a total of nine responses for each instruction.
Step 3: Filtering High-quality Responses. For each instruction, we evaluate nine generated responses using the three models previously mentioned, assigning quality scores to each response. The scores from the three models are then averaged to identify the response with the highest overall score.
Combining the optimal response with the corresponding instruction forms the instruction dataset. The exact prompt we use for scoring is provided in Table 3.
3.2 Data mixing and training.
The reconstructed SFT data is blended with domain-specific SFT data at a controlled ratio, and the combined dataset is used for full fine-tuning.
4 Experiments
4.1 Setup
We evaluate the performance of the general-purpose base LLM after domain-specific fine-tuning in the medical field. Our method-generated dataset is mixed with medical data for SFT, and performance is assessed on common general benchmarks. Comparisons are made against models fine-tuned using open-source third-party SFT datasets.
Base LLM Our experiment used Llama3-70B-Instruction as the base model. Llama3-70B-Instruction has undergone previous instruction fine-tuning during pre-train to improve its instruction following ability.
Baseline Datasets for SFT AWe adopted several common open-source SFT datasets instead of the SFT dataset established by our method mixed with the specific medical datasets as baselines. These datasets include: Evol Instruct (Xu et al., 2023), UltraChat (Ding et al., 2023), ShareGPT (Chiang et al., 2023), WildChat (Zhao et al., 2024),GenQA (Chen et al., 2024), OpenHermes 1 (Teknium, 2023b), OpenHermes 2.5 (Teknium, 2023a), and Tulu V2 Mix (Ivison et al., 2023). ShareGPT and WildChat are representative human-written datasets containing 112K and 652K high-quality multi-round conversations between humans and GPT, respectively. Evol Instruct, UltraChat, and GenQA are representative open-source synthetic datasets. Following Meng et al. (2024), we use the 208K sanitized version of Ultrachat provided by HuggingFace [^1] https://huggingface.co/datasets/HuggingFaceH4/ultrachat_200k. OpenHermes 1, OpenHermes 2.5, and Tulu V2 Mix are crowd-sourced datasets consisting of a mix of diverse open-source instruction datasets, with 243K, 1M, and 326K conversations, respectively.
Evaluation Performance is measured on general-purpose tasks from the Hugging Face Open LLM Leaderboard (Beeching et al., 2023), including MMLU-PRO (general knowledge) (Wang et al., 2024), GPQA (expert-level Q&A) (Rein et al., 2023), IFEval (instruction following) (Zhou et al., 2023) and MATH level 5 (mathematical reasoning) (Hendrycks et al., 2021).
Implementation Details All datasets are standardized to the same total size, with medical data fixed at 17%. Training is conducted via LLaMA-Factory with a learning rate of 1e-6 and 1 epoch on 8 Tesla A800-80GB GPUs.
4.2 Results
As demonstrated in Table 1, the model fine-tuned with our method achieves marked improvements across all four general benchmarks compared to models trained on mixtures of third-party open-source datasets and domain-specific data.
Figure 1: SFT Data Distribution of Base LLM
Alignment Setup MMLU-PRO (5) GPQA (0) IFEval(0) Math Lvl 5 (4) Average
Llama-3-70B-Instruct 46.74 4.92 80.99 23.34 39.00
ShareGPT 46.14 4.31 81.31 20.24 38.00
Evol Instruct 45.76 4.64 82.52 22.30 38.81
GenQA 43.33 4.48 80.43 15.41 35.91
OpenHermes 1 45.31 4.21 81.91 15.52 36.74
OpenHermes 2.5 45.63 4.79 82.33 15.62 37.09
Tulu V2 Mix 46.47 4.19 82.69 16.62 37.49
WildChat 45.83 4.12 81.32 22.11 38.35
UltraChat 45.15 4.08 81.57 20.31 37.78
Ours 46.73 4.88 81.93 23.29 39.21
Table 1: Final result of the SFT model using our method and the official instruction model across various downstream benchmarks.
Alignment Setup MMLU-PRO (5) GPQA (0) IFEval(0) Math Lvl 5 (4) Average
Llama-3-70B-Instruct 46.74 4.92 80.99 23.34 39.00
Single model-Single response 46.11 4.72 81.31 23.11 38.81
Single model-Multi-responses-Filtered 46.55 4.91 81.72 23.06 39.06
Three Models-Multi-responses-Filtered 46.73 4.88 81.93 23.29 39.21
Table 2: Ablation result of our method.
Impact of SFT Data Distribution As illustrated in Figure 1, the distribution of SFT instruction types extracted from the original LLM primarily comprises information-seeking and advice-seeking data, with part proportions of coding data, a small amount of mathematical and reasoning tasks, and other categories. Our reconstructed general dataset, directly generated from the base model’s instructions, strictly adheres to this intrinsic distribution. In contrast, existing open-source SFT datasets exhibit more significant divergence from the original distribution, particularly in their underrepresentation of coding and reasoning tasks, and overrepresentation of generic conversational templates. This misalignment exacerbates catastrophic forgetting when third-party datasets are used for fine-tuning, as the model’s parameters are forced to adapt to a distorted data landscape that conflicts with its pre-training and initial SFT phases.
4.3 Ablation Study
To further examine the impact of individual components in our proposed method, we conduct an ablation study focusing on multi-response generation and filtering high-quality responses.We respectively test another two configurations:
Single model-Single response After instruction generation, only the target model (LlamA3-70B-Instruct) generates a single response, bypassing multi-response generation and filtering.
Single model-Multi-responses-Filtered The target model generates three responses without external models, followed by filtering.
Results are presented in table 2. The single-response setup exhibits performance degradation compared to the full method but still surpasses models fine-tuned with open-source dataset mixtures in table 1 with the three-response setup (all generated by one LLM without external models and filtered) albeit slightly underperforming the complete method.
These findings highlight the critical role of all three stages in our framework. Instruction Generation serves as the foundation, ensuring the data distribution approximates the original. Subsequently, Multi-Response Generation and Filtering High-Quality Responses iteratively refine data quality by synthesizing multiple candidate responses and selecting optimal outputs. Crucially, integrating more capable LLMs (GPT-4 and Qwen2.5-72B-Instruct) as response generators introduces a knowledge distillation effect, where the superior reasoning and generation capabilities of these models enhance the quality of synthesized SFT data. This hierarchical design—first recovering the latent instruction distribution, then refining responses through multi-model collaboration—ensures both distributional fidelity and high-quality supervision signals, which are essential for mitigating catastrophic forgetting while improving task-specific performance.
5 Conclusion
In this paper, we developed a method to extract instruction distributions from a model trained on an unpublished instruction dataset. We then leveraged two additional powerful models to collaboratively generate high-quality responses, forming a dataset used for SFT. Compared to other baseline open-source SFT datasets, our approach mitigates more catastrophic forgetting and enhances general fine-tuning performance.
6 Declaration
Our method is designed to reconstruct an approximate original SFT instruction distribution from the model for enhance SFT performance, not aiming and unable to extract exact training data. Therefore, there is no concern of proprietary data replication.This ensures compliance with ethical and legal standards.
Limitations
We conducted experiments only on Llama-3-70B-Instruct, achieving favorable results. Due to computational constraints, we did not perform extensive testing on other size models and datasets. Furthermore, future work should investigate strategies to seamlessly integrate our approach with existing continual learning methods specifically designed to mitigate catastrophic forgetting. This integration could potentially harness complementary strengths—leveraging our data distribution reconstruction to preserve generalization while employing architectural or regularization-based techniques to stabilize parameter updates—to achieve more robust and comprehensive solutions against knowledge erosion in sequential fine-tuning scenarios.
Appendix A A prompt used to evaluate the quality of a response.
Below is a user instruction and an AI response. Evaluate the quality of the AI’s response based on how well it fulfills the user’s request. Assign a score based on the following 5-point scale: 1: The response is incomplete, off-topic, or contains irrelevant, vague, or missing information. It may repeat the user’s question, include personal opinions, or be written from a non-AI perspective (e.g., blog-like). It may also have promotional or irrelevant content. 2: The response addresses some of the user’s request but lacks detail or direct relevance. It provides only a general approach instead of a specific solution. 3: The response is helpful but lacks an AI perspective. It covers the user’s request but appears taken from a personal blog, webpage, or similar source. It may include personal opinions, experiences, or mentions of external content. 4: The response is clear, complete, and written from an AI’s perspective. It directly addresses the user’s request, but there may be minor room for improvement, such as clarity or conciseness. 5: The response is excellent, written from an AI’s perspective, with a clear focus on the user’s request. It is thorough, well-organized, and shows expert knowledge without irrelevant content. The response is logical, easy to follow, and engaging. Provide a brief justification for your score and then write "Score: <rating>" in the last line. <generated instruction> <output>
Table 3: A prompt used to evaluate the quality of a response.